OpenStudy (anonymous):

During the month of February, Fabulous Feet Shoe Mart sold 48 pairs of red loafers. After an ad campaign to boost sales, they sold 60 pairs in March. Find the percent of increase in sales. A. 29% B. 19% C. 25% D. 15%

OpenStudy (anonymous):

@jigglypuff314

jigglypuff314 (jigglypuff314):

48x = 60 x = 60/48

OpenStudy (anonymous):

D

jigglypuff314 (jigglypuff314):

x = 1.25 -> 125% so from 100% to 125% difference is and increase of 25%

OpenStudy (anonymous):

oh okay, got it :)

jigglypuff314 (jigglypuff314):

area of triangle = (1/2)(base)(height) plug in so 60 = (1/2)(12) (height) so height = 60 / 6 =

OpenStudy (anonymous):

D :)

jigglypuff314 (jigglypuff314):

yep :)

OpenStudy (anonymous):

In the diagram, the figures are similar. What is x?

jigglypuff314 (jigglypuff314):

\[\frac{ x }{ 8 } = \frac{ 3 }{ 7 }\]

OpenStudy (anonymous):

c

jigglypuff314 (jigglypuff314):

not quite x = 8 * 3 / 7 = 3.43

OpenStudy (anonymous):

ok thank you again

OpenStudy (anonymous):

A tree casts a shadow 10 ft long. A boy standing next to the tree casts a shadow 2.5 ft long. The triangle shown for the tree and its shadow is similar to the triangle shown for the boy and his shadow. If the boy is 5 ft tall, how tall is the tree?

jigglypuff314 (jigglypuff314):

\[\frac{ 5 }{ 2.5 } = \frac{ x }{ 10 }\]

OpenStudy (anonymous):

5

OpenStudy (anonymous):

20

jigglypuff314 (jigglypuff314):

20 is correct :)

OpenStudy (anonymous):

thank you so much :-)

OpenStudy (anonymous):

What is the solution of the equation? 4.7x + 3.8 = 13.2 A. 3 B. 3.7 C. 2 D. 3.6

jigglypuff314 (jigglypuff314):

subtract 3.8 from both sides divide both sides by 4.7

OpenStudy (anonymous):

C

jigglypuff314 (jigglypuff314):

that is correct :)

OpenStudy (anonymous):

What is the total cost of a $56.53 meal at a restaurant after including a 17% tip? A. $3.33 B. $9.61 C. $66.14 D. $39.53

jigglypuff314 (jigglypuff314):

56.53 * 1.17 =

OpenStudy (anonymous):

66.14

OpenStudy (anonymous):

John and 2 friends are going out for pizza for lunch. They split one pizza and 3 large drinks. The pizza cost $12.00. They spend a total of $16.95. Find the cost of one large drink. A. $1.70 B. $1.65 C. $14.48 D. $2.48

jigglypuff314 (jigglypuff314):

brb

OpenStudy (anonymous):

b right

OpenStudy (anonymous):

ok

jigglypuff314 (jigglypuff314):

Hi I'm back ;) and B would be correct :)

OpenStudy (anonymous):

hi, and okay thank u :-)

OpenStudy (anonymous):

What is the solution of the equation?

jigglypuff314 (jigglypuff314):

subtract 27 from both sides multiply both sides by 6 divide both sides by 4

OpenStudy (anonymous):

4

jigglypuff314 (jigglypuff314):

4p/6 + 27 = 39 -27 -27 4p/6 = 12 4p = 72 p = 72 / 4 =

OpenStudy (anonymous):

18

OpenStudy (anonymous):

What is the solution of the equation? 16 = –d + 6 A. 10 B. –10 C. –9 D. –15

jigglypuff314 (jigglypuff314):

subtract 6 from both sides multiply both sides by -1

OpenStudy (anonymous):

b

jigglypuff314 (jigglypuff314):

yep :)

OpenStudy (anonymous):

ok :), last one

OpenStudy (anonymous):

What is the solution of the equation?

jigglypuff314 (jigglypuff314):

multiply both sides by 5 subtract both sides by 6

OpenStudy (anonymous):

c

jigglypuff314 (jigglypuff314):

b + 6 = 50 b = 50 - 6 =

OpenStudy (anonymous):

44

jigglypuff314 (jigglypuff314):

yep :)

jigglypuff314 (jigglypuff314):

d = 48 * 14 / 12

OpenStudy (anonymous):

56

jigglypuff314 (jigglypuff314):

yep :)

OpenStudy (anonymous):

thank you for ur help :) will you be on tomorrow?

jigglypuff314 (jigglypuff314):

kind of... the only reason I'm on right now is because I don't have school today :P so tomorrow maybe, late at night...

OpenStudy (anonymous):

ohh :) ok how about tonight? will u be on

jigglypuff314 (jigglypuff314):

yeah sure :) (unless my parents want me to do SAT practice >,<)

OpenStudy (anonymous):

okay haha, ill message you :) are you good at graphing and Inequalities

jigglypuff314 (jigglypuff314):

yeah, I can do those ;)

OpenStudy (anonymous):

okay :-) ill ttyl then!